Two of the sides in this isosceles right triangle measure 18 and 3 . How do I solve this ?
Answer by Theo(3458)

if it's an isosceles triangle, then 2 of the sides must be equal.

if it's a right isosceles triangle, then the other 2 angles must be 45 degrees.

if 2 of the sides are 3 then the hypotenuse must equal sqrt(18).

3^2 + 3^2 = sqrt(18)^2 becomes 9 + 9 = 18 which would be correct.

this is from the pythagorean theorem which states that a^2 + b^2 = c^2 where c is the hypotenuse.

I suspect that each side is 3 and the hypotenuse is sqrt(18).

otherwise you have no solution based on the information given.